Please help. I am stuck on this problem. I have a partial solution I think but I am not understanding how to do this. Could you please explain what I am supposed to be substituting in this problem and help me find the answer -if independent, inconsistent, or dependent>
x-y=3
3x-2y=3

Answer by scianci(186)

Note that the slopes of these equations are 1 and 3/2. Since they are different, the system is independent. Use a linear combination to solve for one of the variables:
2(x - y = 3) --> 2x - 2y = 6
3x - 2y = 3 --> -3x + 2y = -3
-x = -3
x = 3
(3) - y = 3 [using the 1st. equation]
y = 0
Solution is (3 , 0)